Factorise 2x 2 +7x+5 by using perfect square method please please

Factorise 2x2+7x+5 by using perfect square method please please

Grade:8

1 Answers

Arun
5 years ago
Dear student
 
it can be written as
 
2 (x^2 + 7/2 + 5/2) = 2 (x^2 + 2 *7/4 *x + 49/16 + 5/2 – 49/16)
 
 = 2 (x + 7/4)^2 + 5 – 49/8
 
 = 2 (x + 7/4)^2 – 9/8
 
hence (x + 7/4)^2 = 9/16
 
x + 7/4 = ± ¾
 
hence x = – 1 or – 5/2